# Learning to summarize user information for personalized reinforcement learning from human feedback

### v2

@@ Appendix B PLUS-generated summary examples
Appendix C PLUS in the wild
@@
Having established the applicability of PLUS to the real user dataset, we next investigate two possible user cases of PLUS-generated summaries for downstream personalization: Personalizing LLM-as-judges in Alg. 2 and personalizing LLM response generation in Alg. 3. In both cases, PLUS-generated summaries directly enable personalization of proprietary state-of-the-art models without further finetuning.
@@
Appendix D Personalized LLM-as-Judges
Here, we use an LLM-as-a-judge framework where we prompt GPT4 to select which of the two responses from the PRISM dataset would be preferred, and evaluate its accuracy in predicting the right one with and without conditioning on the PLUS user summary. This allows us to test whether the learned summaries can enable zero-shot personalized preference prediction for proprietary state-of-the-art models. We provide both GPT-4.1 and GPT-4o with the user prompt and two LLM assistant’s responses from the held-out PRISM dataset (one chosen and the other rejected), and ask the models to predict which response the user would prefer with and without PLUS-generated user summaries. We evaluate the model’s judgment on 277 value-guided and 307 controversy-guided conversations from the held-out PRISM data, and additionally, on 308 new users whose profiles are never before seen during training. Table 7 shows that GPT4 makes more accurate predictions about the user’s preference when the summaries are given–especially for GPT4o which makes a 40% improvement. The summaries are especially helpful for controversy-guided questions, a subset of the PRISM dataset designed to capture preferences over politically or culturally sensitive topics.
@@ Appendix D Personalized LLM-as-Judges
Appendix E Response Personalization with PLUS-generated Summaries
@@
Appendix F Additional experiments with the PRISM dataset.
@@
Appendix G Existing RLHF techniques
@@
BTL [2] Reward model Prediction
DPL [23] Reward model Prediction
@@
VPL [18] VAE + Reward model Reconstruction + Prediction
@@
Appendix H Training Details
H.1 Datasets
@@
Figure 5: PPO training curves with the PRISM dataset.
H.2 Hyperparameters
@@
Table 11: PPO training hyperparameters. Our co-adaptation framework in Alg. 1 is built on top of OpenRLHF’s ppo training code [10]. Summarizer $\pi_{\theta}$ is implemented as the PPO actor. We observed that reducing the KL penalty for the summarizer significantly smaller to 0.001 worked well for Pets and PRISM.
@@
Table 12: Reward model training hyperparameters; reward modeling code was adapted from  Poddar et al. [18]. We additionally conducted hyperparameter sweeping for VPL, BTL, DPL, and the oracle reward models by varying the learning rates $\{3*10^{-4},9*10^{-6}\}$ and training batch sizes $\{32,128,256\}$
Appendix I LLM prompts
I.1 Prompt for generating summaries with PLUS.
@@ Appendix I LLM prompts
  • PRISM: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preferences and reflect that understanding in your future responses. ##Past conversation 0## ###Prompt### ###Chosen response### ###Rejected response###… Based on this information, provide a short summary of the user’s preferences to guide how the AI assistant should respond in future conversations. Reflect on the user’s past chosen and rejected responses, and what the user preferred about the chosen responses. Do not focus on the topics of specific conversations. Keep your summary concise–under 50 words. Start your response with ##Summary
@@
I.2 Prompt for oracle reward models.
@@
Appendix J Prompt for In-context Learning Models.
@@
J.1 Prompts for zero-shot personalization with PLUS.
Prompt for LLM-as-a-Judge:
@@
Prompt for generating a new response to user’s query:

### v3

@@ Appendix C PLUS-generated summary examples
Appendix D PLUS in the wild
@@
Having established the applicability of PLUS to the real user dataset, we next investigate two possible user cases of PLUS-generated summaries for downstream personalization: Personalizing LLM-as-judges in Alg. 2 and personalizing LLM response generation in Alg. 3. In both cases, PLUS-generated summaries directly enable personalization of strong proprietary models (e.g., GPT-4) without further finetuning.
@@
Appendix E Personalized LLM-as-Judges
Here, we use an LLM-as-a-judge framework where we prompt GPT4 to select which of the two responses from the PRISM dataset would be preferred, and evaluate its accuracy in predicting the right one with and without conditioning on the PLUS user summary. This allows us to test whether the learned summaries can enable personalized preference prediction for strong, proprietary models without further fine-tuning. We provide both GPT-4.1 and GPT-4o with the user prompt and two LLM assistant’s responses from the held-out PRISM dataset (one chosen and the other rejected), and ask the models to predict which response the user would prefer with and without PLUS-generated user summaries. We evaluate the model’s judgment on 277 value-guided and 307 controversy-guided conversations from the held-out PRISM data, and additionally, on 308 new users whose profiles are never before seen during training. Table 9 shows that GPT4 makes more accurate predictions about the user’s preference when the summaries are given–especially for GPT4o which makes a 40% improvement. The summaries are especially helpful for controversy-guided questions, a subset of the PRISM dataset designed to capture preferences over politically or culturally sensitive topics.
@@ Appendix E Personalized LLM-as-Judges
Appendix F Response Personalization with PLUS-generated Summaries
@@
We selected examples of PLUS-summary conditioned responses generated with GPT-4o and GPT-4.1. The highlighted parts in blue reflect personalization based on the corresponding user summary, especially when compared to the default GPT responses.
@@
Appendix G Additional experiments with the PRISM dataset.
@@
Appendix H Existing RLHF techniques
@@
BTL (Bradley and Terry, 1952) Reward model Prediction
DPL (Siththaranjan et al., 2024) Reward model Prediction
@@
VPL Poddar et al. (2024) VAE + Reward model Reconstruction + Prediction
@@
Appendix I Training Details
I.1 Datasets
@@
I.2 Hyperparameters
@@
Table 13: PPO training hyperparameters. Our co-adaptation framework in Alg. 1 is built on top of OpenRLHF’s ppo training code (Hu et al., 2024). Summarizer $\pi_{\theta}$ is implemented as the PPO actor. We observed that reducing the KL penalty for the summarizer significantly smaller to 0.001 worked better for Pets and PRISM. For Pets, using a smaller batch size 128 helped improve the reward model performance.
@@
Table 14: Reward model training hyperparameters; reward modeling code was adapted from  Poddar et al. (2024). We additionally conducted hyperparameter sweeping for VPL, BTL, DPL, and the oracle reward models by varying the learning rates $\{3*10^{-4},9*10^{-6}\}$ and training batch sizes $\{32,128,256\}$. We observed that the oracle models on the UltraFeedback datasets achieve higher accuracy by training for more than 1 epoch.
Appendix J LLM prompts
J.1 Prompt for generating summaries with PLUS.
@@
  • PRISM: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preferences and reflect that understanding in your future responses. ##Past conversation 0## ###Prompt### ###Chosen response### ###Rejected response###… Based on this information, provide a short summary of the user’s preferences to guide how the AI assistant should respond in future conversations. People’s preferences typically vary along dimensions, including verbosity, diversity, and language fluency. Do not focus on the topics of particular conversations; instead, focus on the preferences revealed through past conversation examples. Keep your summary concise-under 100 words. Start your response with ##Summary
@@
J.2 Prompt for oracle reward models.
@@
Appendix K Prompt for In-context Learning Models.
@@
K.1 Prompts for GPT-4 personalization with PLUS.
Prompt for LLM-as-a-Judge.
@@
Prompt for generating a new response to user’s query.